OpenStudy (anonymous):

Jessica deposits $300 into a savings account that pays an annual interest rate of 2%, compounded twice a year. How much money will Jessica have in her account at the end of one year?

OpenStudy (anonymous):

to get the effective annual interest rate, use this formula: i-e=(1+i/2)^2 - 1 = .0201 or 2.01% so then at the end of the year, Jessica will have 300(1+i-e) = $306.03

OpenStudy (anonymous):

that it or wut

OpenStudy (anonymous):

where does the "e" come from? the formula is P=PRT?

OpenStudy (anonymous):

well, generically, which is what i meant to type: A = P(1+r/n)^nt where P = starting amount, r = rate, n=compounded cycle, and t = years

OpenStudy (anonymous):

basically its 306.03

OpenStudy (anonymous):

its under the k12 answers

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

np medal me

OpenStudy (anonymous):

how? lol